# Towards Precise 3D Quantum Control of a Levitated Dipolar Scatterer using Spatial Mode Decomposition

### v1

@@ VII Temperature Estimation
In Supplementary Material (VI) we claimed that an optically-levitated nanoparticle has a motional power spectral density (PSD) given by equation (S29). From the reciprocity of the Wiener-Khinchin theorem, this implies a mean-square displacement along the $q$-axis of the form
@@
$\langle q(t)q(t)\rangle$ $=\frac{k_{B}T_{0}}{m\left(\Omega_{q}+\delta\Omega_{q}\right)^{2}}\frac{\gamma}{\gamma+\delta\gamma^{(q)}}~{}.$ (S32)
@@
According the equipartition theorem of statistical mechanics, this quantity is related to the center-of-mass (COM) temperature $T_{(q)}$ of the nanoparticle, such that
$T_{q}$ $=T_{0}\frac{\gamma}{\gamma+\delta\gamma^{(q)}}~{}.$ (S33)
@@
Here, we use this relationship to estimate the temperature along each spatial axis after a nanoparticle has undergone parametric feedback cooling. To understand this process, consider fitting equation (S30) to experimental data of $S_{(qq)}(\Omega)$ taken at thermal equilibrium ($T_{0}=300$ K). From this, one recovers fit parameters $A_{0}$, $B_{0}$ and $C_{0}$, respectively. Conversely, for data taken of a cooled nanoparticle, we find fit parameters $A_{\mathrm{fb}}$, $B_{\mathrm{fb}}$ and $C_{\mathrm{fb}}$. Hence, we calculate a COM temperature for the cooled nanoparticle,
$T_{q}$ $=T_{0}\frac{A_{\mathrm{fb}}C_{0}}{C_{\mathrm{fb}}A_{0}}~{}.$ (S34)
VIII Calculation of Minimum Achievable Temperature
@@
Here, we briefly calculate the theoretical performance limits of our setup. In particular, we are interested in the minimum achievable temperature(s). For strong feedback signals, we expect a minimum temperature of  [27]
$T_{q}^{\mathrm{min}}$ $=2\sqrt{T_{0}\gamma\frac{\pi m}{2k_{B}}\Omega_{q}^{2}S_{\mathrm{imp},q}}~{}.$ (S35)
@@
For the results discussed here, where are limited to pressures on the order of $5\times 10^{-5}$ mbar, this implies $(T_{x}^{\mathrm{min}},T_{y}^{\mathrm{min}},T_{z}^{\mathrm{min}})=(29,3.7,2.8)\pm(4,0.6,0.4)$ mK. Note that the pressure dependence of these values is encoded in the natural damping rate, $\gamma$, of the nanoparticle’s motion.

### v2

@@ Appendix H Temperature estimation
In Appendix G we claimed that an optically-levitated nanoparticle has a motional power spectral density (PSD) given by Eq. (46). From the reciprocity of the Wiener-Khinchin theorem, this implies a mean-square displacement along the $q$-axis of the form
@@
$\langle q(t)q(t)\rangle$ $=\frac{k_{B}T_{0}}{m\left(\Omega_{q}+\delta\Omega_{q}\right)^{2}}\frac{\gamma}{\gamma+\delta\gamma_{q}}~{}.$ (49)
@@
According the equipartition theorem of statistical mechanics, this quantity is related to the center-of-mass (COM) temperature $T_{q}$ of the nanoparticle, such that
$T_{q}$ $=T_{0}\frac{\gamma}{\gamma+\delta\gamma_{q}}~{}.$ (50)
@@
Here, we use this relationship to estimate the temperature along each spatial axis after a nanoparticle has undergone parametric feedback cooling. To understand this process, consider fitting Eq. (47) to experimental data of $S_{qq}(\Omega)$ taken at thermal equilibrium ($T_{0}=300$ K). From this, one recovers fit parameters $A_{0}$, $B_{0}$ and $C_{0}$, respectively. Conversely, for data taken of a cooled nanoparticle, we find fit parameters $A_{\mathrm{fb}}$, $B_{\mathrm{fb}}$ and $C_{\mathrm{fb}}$. Hence, we calculate a COM temperature for the cooled nanoparticle,
$T_{q}$ $=T_{0}\frac{A_{\mathrm{fb}}C_{0}}{C_{\mathrm{fb}}A_{0}}~{}.$ (51)
Appendix I Calculation of the minimum achievable temperature
@@
Here, we briefly calculate the theoretical performance limits of our setup. In particular, we are interested in the minimum achievable temperature(s). For strong feedback signals, following [16], we estimate the minimum temperature as
$T_{q}^{\mathrm{min}}$ $=2\sqrt{T_{0}\gamma\frac{\pi m}{2k_{B}}\Omega_{q}^{2}S_{\mathrm{imp},q}^{\ell,m}}~{}.$ (52)
@@
For the results discussed here, where are limited to pressures on the order of $5\times 10^{-5}$ mbar, this implies $(T_{x}^{\mathrm{min}},T_{y}^{\mathrm{min}},T_{z}^{\mathrm{min}})=(3.7,29,2.8)\pm(0.6,4,0.4)$ mK. Note that the pressure dependence of these values is encoded in the natural damping rate, $\gamma$, of the nanoparticle’s motion.
Appendix J Influence of polarisation on demultiplexing performance
Figure 6: The demultiplexing observed in each output channel of the spatial mode sorter as the polarisation ($\theta$) of the trapping beam is varied using a half waveplate (HWP), leading to modifications in the electric field profile of the inelastically scattered field components. Here, $\theta=0^{\mathrm{o}}$ refers to a horizontally polarised trapping beam ($\hat{\bm{e}}_{x}$) and $\theta=45^{\mathrm{o}}$ to a vertically polarised trapping beam polarised ($\hat{\bm{e}}_{y}$), respectively.
Here we show the effect that varying the polarisation of the trapping beam has on the demultiplexing performance. We find that there are specific linear polarisations where optimal demultiplexing is observed. That is, where each channel is primarily coupled to the inelastically scattered electric field generated by a unique translational degree of freedom $q(t)$ for $q=x,y$ and $z$. From the experimental results shown below, we find that the optimal case occurs for horizontally ($\hat{\bm{e}}_{x}$) and vertically ($\hat{\bm{e}}_{y}$) polarised trapping beams. Interestingly, we find that depending on the choice of polarisation, $x(t)$ and $y(t)$ are detected in alternating LP11a/b channels. Conversely, the LP01 channel is consistently coupled to $z(t)$, with only minor contributions from the spectrally adjacent mode. This is shown in Fig 6 below, and agrees with those results shown in Fig. 2(a) of the main text.